OpenStudy (anonymous):

What is the radius of the following circle? x^2+y^2=15

OpenStudy (amistre64):

it might help to recall the general equation of a circle, then its just a matter of comparing parts

OpenStudy (anonymous):

Whats the equation??

OpenStudy (amistre64):

it should be in your study material, if not, its posted online in about infinity+1 places ... as a memory device, it is pythagorean

sam (.sam.):

The general equation is \[x^2+y^2=r^2\] r is the radius, Then you can see that \[r^2=15\] Then \[r=?\]